# A comparison of translation performance between DeepL and Supertext

### v1

@@ Abstract
As strong machine translation (MT) systems are increasingly based on large language models (LLMs), reliable quality benchmarking requires methods that capture their ability to leverage extended context. This study compares two commercial MT systems – DeepL and Supertext – by assessing their performance on unsegmented texts. We evaluate translation quality across four language directions with professional translators assessing segments with full document-level context. While segment-level assessments indicate no strong preference between the systems in most cases, document-level analysis reveals a preference for Supertext in three out of four language directions, suggesting superior consistency across longer texts. We advocate for more context-sensitive evaluation methodologies to ensure that MT quality assessments reflect real-world usability. [^1] We release all evaluation data and scripts for further analysis and reproduction at https://github.com/supertext/evaluation_deepl_supertext
@@ 1 Introduction
After the transition from statistical to neural modelling roughly a decade ago (Kalchbrenner and Blunsom, 2013; Sutskever et al., 2014; Bahdanau et al., 2015), the field of MT is undergoing another paradigm shift towards leveraging LLMs Xu et al. (2024); Wu et al. (2024b); Kocmi et al. (2024). LLM-based translation offers the potential for significantly improved translation quality, especially with respect to consistent translation of documents. Unlike neural machine translation (NMT) systems, which typically process documents as isolated sentences or paragraphs (Post and Junczys-Dowmunt, 2023), many LLMs operate with context windows that can span thousands of words, allowing them to maintain consistency throughout a document – for instance, by ensuring that a word’s translation in the final sentence aligns with its rendering in the first Wu et al. (2024b).
In the most recent shared task at the Conference on Machine Translation (WMT24) that focuses on evaluating the state of the art in general-domain translation quality, the majority of the 28 system submissions were already based on LLMs Kocmi et al. (2024). Although without statistical significance and for the language direction English to German only, one system even outranked the human reference translations as evaluated by professional human annotators.
Despite this impressive achievement, findings of human-machine parity should be approached with caution. Similar claims already emerged with pre-LLM technology Hassan et al. (2018), yet have subsequently been refuted due to limitations in the evaluation design focusing on single segments in isolation Läubli et al. (2018); Toral et al. (2018); Freitag et al. (2021).
The WMT24 shared task also highlights that evaluations based on automatic metrics (rather than human evaluation) can lead to wrong conclusions when comparing strong MT systems Kocmi et al. (2024).
@@
In this paper, we evaluate two commercial translation systems (Section 2) under conditions that allow for leveraging the full-text capabilities of LLMs. Source texts are translated without prior segmentation (Section 3), and the resulting translations are rated by professional translators considering the full document context (Section 4). We find that while both systems translate a similar number of segments better than the other, the difference is more pronounced on the document level (Section 5), which we attribute to differences in how much context the systems consider during translation (Section 6). Our findings suggest that the adoption of LLMs creates opportunities for smaller players to challenge dominant industry leaders (Section 7).
@@ 2 Systems
DeepL [^3] https://www.deepl.com is a widely used MT provider boasting ‘unrivalled translations that set the standard’. [^4] https://www.deepl.com/en/quality In the latest Intento report, it scores best among nine ‘real-time engines’ and, together with GPT-4, is found to ‘consistently outperform other models’.2 Due to its closed source, the technology behind DeepL’s translation system is not publicly known.
@@
While both DeepL and Supertext provide target language variants for English (British and American), Supertext also provides target language variants for German (Austria, Germany, Switzerland), French (France, Switzerland), and Italian (Italy, Switzerland). We use a specific target language variant whenever available (Section 3.2).
@@ 3.1 Source Texts
We collect 20 texts each in two source languages: English (en) and German (de). All texts stem from news websites: New York Times [^6] https://www.nytimes.com for English and Neue Zürcher Zeitung [^7] https://www.nzz.ch for German, respectively. We select 10 FAQ pages and 10 recent news articles in the economy section from each website. To balance the distribution of text lengths, we cropped the end of some documents omitting their final paragraphs. Notably, these texts are only available in a single language; they are unlikely to be contained in the training data of either system we evaluate.
@@ 3.2 Target Texts
We create translations in four language directions (Table 1) directly in the respective online translation interface of each system as a regular user would. [^8] All translations were produced on 27 January 2025. We do not modify the texts before translation and paste them in their original formatting, including newlines.
After translation, we manually split and align the source texts and translations into sentences. If one of the systems merges two or more sentences into a single sentence, we ensure that the same content is merged for the other system, such that the raters are presented with parallel segments. Table 1 shows the number of segments per language pair.
Language Direction Texts Segments
de $\rightarrow$ en-GB 20 281
de $\rightarrow$ fr-CH 20 276
de $\rightarrow$ it-CH 20 265
en $\rightarrow$ de-CH 20 211
@@ 4.2 Materials
We arrange all segments of a source document with their corresponding translations by both systems in a spreadsheet. The segments are presented in original document order, including formatting such as newlines, such that raters see the full source text and both translations side-by-side. We randomly assign the system outputs to columns labelled Translation A and Translation B for each text such that raters do not have any information about which translation stems from which system (a blind A/B test setting). Assignments are kept consistent within a text such that the document context remains natural.
@@ 7 Conclusion
While A/B tests are commonly used for comparing two systems and a reliable basis for incrementally improving MT systems Tang et al. (2010); Wu et al. (2024a), they provide no insight into the severity of errors within a translation or across different systems compared to MQM ratings (Freitag et al., 2021). Moreover, some mistakes may be harder to spot than others during real-world usage when not being shown contrastively against an alternative translation. Similarly, the preference in an A/B test may not correlate with the effort needed for post-editing the translation. To address these questions, we plan to extend our evaluation efforts.
@@
Finally, the scope of this study is not exhaustive but is limited to a subset of language pairs, two domains, and a limited number of documents. Yet, we are providing details that go beyond what DeepL is sharing publicly.4

### v3

@@ Abstract
As strong machine translation (MT) systems are increasingly based on large language models (LLMs), reliable quality benchmarking requires methods that capture their ability to leverage extended context. This study compares two commercial MT systems – DeepL and Supertext – by assessing their performance on unsegmented texts. We evaluate translation quality across four language directions with professional translators assessing segments with full document-level context. While segment-level assessments indicate no strong preference between the systems in most cases, document-level analysis reveals a preference for Supertext in three out of four language directions, suggesting superior consistency across longer texts. We advocate for more context-sensitive evaluation methodologies to ensure that MT quality assessments reflect real-world usability. [^1] Paper accepted at MT Summit 2025.222We release all evaluation data and scripts for further analysis and reproduction at https://github.com/supertext/evaluation_deepl_supertext
@@
Alex Flückiger, Chantal Amrhein, Tim Graf, Frédéric Odermatt, Martin Pömsl, Philippe Schläpfer, Florian Schottmann, Samuel Läubli Supertext {firstname.lastname}@supertext.com
footnotetext: ©  2025 The authors. This article is licensed under a Creative Commons 4.0 licence, no derivative works, attribution, $C$-BY-ND.
@@ 1 Introduction
After the transition from statistical to neural modelling roughly a decade ago (Kalchbrenner and Blunsom, 2013; Sutskever et al., 2014; Bahdanau et al., 2015), the field of MT is undergoing another paradigm shift towards leveraging LLMs (Xu et al., 2024; Wu et al., 2024b; Kocmi et al., 2024). LLM-based translation offers the potential for significantly improved translation quality, especially with respect to consistent translation of documents. Unlike neural machine translation (NMT) systems, which typically process documents as isolated sentences or paragraphs (Post and Junczys-Dowmunt, 2023), many LLMs operate with context windows that can span thousands of words, allowing them to maintain consistency throughout a document – for instance, by ensuring that a word’s translation in the final sentence matches its previous forms (Wu et al., 2024b).
In the most recent shared task at the Conference on Machine Translation (WMT24) that focuses on evaluating the state of the art in general-domain translation quality, the majority of the 28 system submissions were already based on LLMs (Kocmi et al., 2024). Although without statistical significance and for the language direction English to German only, one system even outranked the human reference translations as evaluated by professional human annotators.
Despite this impressive achievement, findings of human-machine parity should be approached with caution. Similar claims already emerged with pre-LLM technology (Hassan et al., 2018), yet have subsequently been refuted due to limitations in the evaluation design focusing on single segments in isolation (Läubli et al., 2018; Toral et al., 2018; Freitag et al., 2021). The WMT24 shared task also highlights that evaluations based on automatic metrics (rather than human evaluation) can lead to wrong conclusions when comparing strong MT systems (Kocmi et al., 2024).
@@
In this paper, we evaluate two commercial translation systems (Section 2) under conditions that allow for leveraging the full-text capabilities of LLMs. The segmentation of the source text is handled by the translation systems alone without any prior splitting (Section 3), and the resulting translations are rated by professional translators considering the full document context (Section 4). We find that while both systems translate a similar number of segments better than the other, the difference is more pronounced on the document level (Section 5), which we attribute to differences in how much context the systems consider during translation (Section 6). Our findings suggest that the adoption of LLMs creates opportunities for smaller players to challenge dominant industry leaders (Section 7).
@@ 2 Systems
DeepL [^4] https://www.deepl.com is a widely used MT provider boasting ‘unrivalled translations that set the standard’. [^5] https://www.deepl.com/en/quality, see also Appendix A. In the latest Intento report, it scores best among nine ‘real-time engines’ and, together with GPT-4, is found to ‘consistently outperform other models’.3 Due to its closed source, the technology behind DeepL’s translation system is not publicly known.
@@
While both DeepL and Supertext provide target language variants for English (British and American), Supertext also provides target language variants for German (Austria, Germany, Switzerland), French (France, Switzerland), and Italian (Italy, Switzerland). As our use case is machine translation for people in Switzerland, we use the Swiss target language variant whenever available (Section 3.2). [^7] Compared to English variants, the Swiss variants of other languages differ minimally.
@@ 3.1 Source Texts
We collect 20 texts each in two source languages: English (en) and German (de). All texts stem from news websites: New York Times [^8] https://www.nytimes.com for English and Neue Zürcher Zeitung [^9] https://www.nzz.ch for German, respectively. We select 10 FAQ pages and 10 recent news articles in the economy section from each website. Notably, these texts are only available in a single language; they are unlikely to be contained in the training data of either system we evaluate. To balance the distribution of text lengths, we trim the end of some texts by omitting their final paragraphs.
@@ 3.2 Target Texts
We create translations in four language directions (Table 1) directly in the respective online translation interface of each system as a regular user would. [^10] All translations were produced on 27 January 2025. We do not modify the texts before translation and paste them in their original formatting, including newlines. The translation systems may segment the text into smaller chunks internally.
After translation, we manually split and align the source texts and translations into sentences. If one of the systems merges two or more sentences into a single sentence, we ensure that the same content is merged for the other system, such that the raters are presented with parallel segments. Table 1 shows the resulting number of segments per language pair. The texts per source language are identical, differing only in how they were manually segmented for the A/B test after translation. Across the language pairs, the median number of segments per document is 13.
Language Direction Texts Segments Words
de $\rightarrow$ en-GB 20 281 3336
de $\rightarrow$ fr-CH 20 276 3336
de $\rightarrow$ it-CH 20 265 3336
en $\rightarrow$ de-CH 20 211 3483
Total 80 1033 13491
@@ 4.2 Materials
We arrange all segments of a source document with their corresponding translations by both systems in a spreadsheet. The segments are presented in original document order, including formatting such as newlines, such that raters see the full source text and both translations side-by-side. We randomly assign the system outputs to columns labelled Translation A and Translation B for each text such that raters do not have any information about which translation stems from which system (a blind A/B test setting). System assignments are kept consistent within a text such that the document context remains natural.
@@ 7 Conclusion
While A/B tests are commonly used for comparing two systems and a reliable basis for incrementally improving MT systems (Tang et al., 2010; Wu et al., 2024a), they provide no insight into the severity of errors within a translation or across different systems compared to MQM ratings (Freitag et al., 2021). Absent the use of more time-intensive evaluation frameworks, such limitations persist irrespective of whether preferences are aggregated at the system level or pooled by document.
During real-world usage, some mistakes may be harder to spot than others when not being shown contrastively against an alternative translation. Similarly, the preference in an A/B test may not correlate with the effort needed for post-editing the translation. To address these questions, we plan to extend our evaluation efforts.
@@
Finally, the scope of this study is not exhaustive but is limited to a subset of language pairs, two domains, and a limited number of documents. Yet, we are providing details that go beyond what DeepL is sharing publicly on their website.5
@@ Acknowledgments
Appendix A Appendix
For the sake of persistency, we share the archived link as well: https://web.archive.org/web/20250215011944/https://www.deepl.com/en/quality